Grade 11General Physics

If energy is E velocity V and time T are taken as fundamental quantities.calculate dimensions of surface tension

aashu
9 Years agoGrade 11

1 Answer

Ashmak Moon
9 Years ago
SI unit of tension is N/m, we have to write its dimensions in terms of [EVT], which is [(Nm), (m/s), (s)]. Hence we can write N/m in terms of EVT as [E^a  V^b  T^c] where a,b and c are some variables. We put the the SI expansions for EVT as [(Nm)^a  (m/s)^b  (s)^c]and write equations in a,b and c to get N/m as a=1, a+b= (-1) and (-b)+c=0 to give a= 1, b= (-2) and c= (-2). Hence [T] = [E  V^-2  T^-2]